# Local Level Dynamic Random Partition Models for Changepoint Detection

### v1

@@ 6 Discussion
We introduced a new approach for modeling partitions with temporal dependence that relies on the definition of partition-based state equations in a state-space model, thus extending the Dynamic Linear Model widely used in multivariate time series analysis. By characterizing the evolution of the local level equation across partitions and detecting changepoints, our model effectively captures changes in cluster structures over time. Our approach maximizes the use of temporal information while introducing dependence only when supported by the data. This is achieved by incorporating a spike-and-slab formulation to model independence between partitions, thereby enhancing model flexibility. We have illustrated the model’s performance using synthetic data, which hihglighted its accuracy in recovering cluster structures and correctly identifying changepoints. In an application to human gesture data, our model provides reasonable results, particularly in identifying most of the changepoints during activity phases. These results offer valuable insights into the mechanics of human movement and have the potential to significantly enhance gesture recognition systems. Furthermore, the partitions estimated by our random partition model could be integrated as predictors into broader outcome models. For instance, they could be used to determine how different gesture intervals are associated with a propensity for risk-taking or other behavioral outcomes. Thus, understanding these gesture phase clusters can lead to a deeper insight into human behavior, accounting for all probabilistic uncertainties.
Drawing upon the framework presented in Section 2.2, our model can be extended to analyze multivariate data across multiple sources, where dependences arise beyond temporal factors, e.g., varying experimental conditions. This adaptation allows our model to accommodate diverse data structures and complex inter-condition dependences. Importantly, even in this extended context, the random partitions maintain a marginal identical distribution, e.g. the distribution of a Chinese restaurant process. This underscores the model’s adaptability and versatility in capturing complex dependences across different datasets.
@@
Finally, an exciting avenue for future research is the development of dependent random partition state-space models capable of modeling latent state equations that encode more complex dynamics. This extension could encompass higher-order time dependences, allowing for exploring more complex temporal relationships and patterns in data. In order to ensure smoother transitions over time, a hierarchical model formulation could shrink the base partition distribution $p^{*}(\pi)$ to the partition estimated at time $t-1$ or a fixed “anchor” partition. However, care will be needed to ensure computational feasibility and interprability of the parameters $\gamma$. See, e.g., Paganin et al. (2021) or Dahl et al. (2023) for details on recently proposed anchored partitioned models. Similarly, we have assumed that the change point selection indicator $\gamma_{i,t}$ is independent across times. More in general, the changepoint selection could depend on time-varying covariates, such as respiratory data or measurements of expended effort in the analysis of human gesture data. Such developments have the potential to significantly enhance our ability to capture and understand time-varying clustering structures in complex time series data across a wide range of domains and applications.
@@
The Supplementary Material is organized as follows. In Section A we provide the proofs of the results in Section 2 of the main article. Section B provides additional details on posterior computations. Finally, Sections C and D report further information on the simulation studies and the application, in Sections 4 and 5 of the main article.
@@ A.1 Proof of Proposition 1
We let $c_{i,t}$, for $i=1,\ldots,n$ and $t=1,\ldots,T$, be a random variable taking values in $\{1,\ldots,|\pi_{t}|\}$, such that $c_{i,t}=j$ if $i\in C_{j,t}$. As observed in Page et al. (2022), the ERI can be written as
$\varphi_{t_{1},t_{2}}=\mathbbm{E}[R(\pi_{t_{1}},\pi_{t_{2}})]=\binom{n}{2}^{-1}\sum_{1\leq i\leq j\leq n}\psi_{i,j},$
where $\psi_{i,j}=P(c_{i,t_{1}}=c_{j,t_{1}},c_{i,t_{2}}=c_{j,t_{2}})+P(c_{i,t_{1}}\neq c_{j,t_{1}},c_{i,t_{2}}\neq c_{j,t_{2}})$. Following exchangeability, it is easy to verify that $\varphi_{t_{1},t_{2}}=\psi_{1,2}$. From the predictive distribution of a Gibbs-type prior, we obtain $P(c_{1,t_{1}}=c_{2,t_{1}})=V_{2,1}(1-\sigma)$ and $P(c_{1,t_{1}}\neq c_{2,t_{1}})=V_{2,2}$. Moreover,
@@ B.1 Estimating the marginal likelihood $g_{t}$
    If the kernel $p(y_{i,t}\mid\beta_{i,t})$ in (2), and the base measure $P_{0}(\cdot)$ are specified in a convenient form, the function $p(\bm{Y}_{t}\mid\pi_{t})$ is available in closed-form and easy to evaluate. For example, when both are assumed normal (specifically $y_{i,t}\mid\beta_{i,t}\sim\text{N}(\beta_{i,t},\tau^{2})$ and $P_{0}(\cdot)$ Normal with mean $\mu$ and variance $\varsigma^{2}$), as implemented in Sections 4 and 5, then
@@ B.2 Reshuffling step
Updating the partitions $\bm{\pi}_{1:T}$, conditionally on $\bm{\gamma}_{2:T}$, at each iteration of the Gibbs sampler helps improving the mixing of the algorithm. We discuss how this can be done efficiently. We start by observing that the values taken by $\bm{\gamma}_{2:T}$ determine a partition of the partitions $\pi_{1:T}$ into $k_{T}\leq T$ blocks. Within each block, all partitions are identical. We introduce the notation $\{\pi_{1}^{*},\ldots,\pi_{k_{T}}^{*}\}$ to indicate the $k_{T}$ partitions corresponding to the $k_{T}$ blocks and we observe that there can be ties among the elements of $\{\pi_{1}^{*},\ldots,\pi_{k_{T}}^{*}\}$. For example, if $\gamma_{2}=\gamma_{3}=0$ while $\gamma_{4}=1$, then the first block of partitions consists of $\pi_{1}^{*}=\pi_{1}=\pi_{2}=\pi_{3}$. The reshuffling step thus consists in updating the values $\pi_{j}^{*}$, for $j=1,\ldots,k_{T}$, from the corresponding full conditional distributions. If $\mathcal{D}_{j}$ denotes the set of partitions assigned to the $j$th cluster, then the full conditional of $\pi_{j}^{*}$ takes the form
@@
If $|\mathcal{D}_{j}|=1$, say $\mathcal{D}_{j}=\{\pi_{t_{j}}\}$, for some $t_{j}$, then the problem of sampling from (19) can be tackled simply by picking an element at random from the list $\mathscr{S}_{t_{j}}$. If $|\mathcal{D}_{j}|>1$, say $\mathcal{D}_{j}=\{\pi_{t_{j,1}},\ldots,\pi_{t_{j,|\mathcal{D}_{j}|}}\}$, then we propose a sampling importance resampling step to simulate from (19). Specifically,
@@
    For every $r=1,\ldots,|\mathcal{D}_{j}|$, we sample at random $m_{t_{j,r}}$ draws from $\mathscr{S}_{t_{j,r}}$ (i.e. from the conditional distribution of $\pi_{t_{j,r}}$ given $\bm{Y}_{t_{j,r}}$). In total we thus get $m_{j}=m_{t_{j,1}}+\ldots+m_{t_{j,|\mathcal{D}_{j}|}}$ candidate partitions from $\mathscr{S}_{j}^{\ast}=\bigcup_{r\in\mathcal{D}_{j}}\mathscr{S}_{t_{j,r}}$.
@@
    We assign a weight $w_{\ell}$ to each partition $\tilde{\pi}_{\ell}$ in $\mathscr{S}_{j}^{\ast}$, with $\ell=1,\ldots,m_{j}$. Each weight is computed, up to a proportionality constant, by evaluating (19) at $\pi_{j}^{\ast}=\tilde{\pi}_{\ell}$.
@@ Appendix C More on the simulation studies of Section 4
We present additional plots related to the simulated data in Section 4.

### v2

@@ 6 Discussion
We introduced a new approach for modeling partitions with temporal dependence that relies on the definition of partition-based state equations in a state-space model, thus extending the dynamic linear model widely used in multivariate time series analysis. By characterizing the evolution of the local level equation across partitions and detecting changepoints, our model effectively captures changes in cluster structures over time. Our approach maximizes the use of temporal information while introducing dependence only when supported by the data. This is achieved by incorporating a mixture representation, reminiscent of the spike-and-slab prior commonly used in variable selection, to model the dependence between partitions. Here, a changepoint is defined as a shift in the partition of units between two consecutive time points. If this shift in cluster allocations coincides with a change in the process mean, our model can detect the changepoint in the traditional sense. However, if the process mean changes while the partition of units remains unchanged, the model will not identify a changepoint. This behavior reflects the model’s design: it is tailored to detect changes in clustering structure rather than in the underlying process itself. We have illustrated the model’s performance using synthetic data, which highlighted its accuracy in recovering cluster structures and identifying changepoints. In an application to human gesture data, our model yields interpretable results by identifying most changepoints during periods of activity. These results offer valuable insights into the mechanics of human movement and have the potential to significantly enhance gesture recognition systems. Furthermore, the partitions estimated by our random partition model could be integrated as predictors into broader outcome models. For instance, they could be used to determine how different gesture intervals are associated with a propensity for risk-taking or other behavioral outcomes. In this context, identifying gesture phase clusters while fully accounting for probabilistic uncertainty may offer a deeper understanding of human behavior.
Drawing upon the framework presented in Section 2.2, our model can be extended to analyze multivariate data across multiple sources, where dependences arise beyond temporal factors, e.g., varying experimental conditions. This adaptation allows our model to accommodate diverse data structures and complex inter-condition dependences. Importantly, even in this extended context, the random partitions maintain a marginal identical distribution, e.g. the distribution of a CRP. Although extending our algorithm for partition changepoint detection to multiview scenarios is not straightforward, an application to epidemiological data displays its effectiveness in a two-view clustering context. Another potential extension involves applications where multiple variables are observed across multiple subjects. Our model can be adapted to accommodate this added complexity, provided the data follow an experimental design in which changepoints are expected to be shared across subjects and the subjects are exchangeable. This structure does not apply to the human gesture dataset described in Section 5.1, where, although data are available for multiple individuals, each subject recounts a different comic story. As a result, the time series reflect distinct underlying processes, and changepoints are not shared across subjects. A related example is offered by Ricci et al. (2024), who recently proposed a biclustering approach that groups subjects based on similarities in temporal trends observed during a task-based fMRI experiment.
@@
Another exciting avenue for future research is the development of dependent random partition state-space models capable of modeling latent state equations that encode more complex dynamics. This extension could encompass higher-order time dependences, allowing for exploring more complex temporal relationships and patterns in data. In order to ensure smoother transitions over time, a hierarchical model formulation could shrink the base partition distribution $p^{*}(\cdot)$ to the partition estimated at the previous time or to a fixed “anchor” partition. However, care will be needed to ensure computational feasibility and interpretability of the latent variables $\gamma_{t}$. See, e.g., Paganin et al. (2021) or Dahl et al. (2023) for details on recently proposed anchored partitioned models. On a related note, while we have assumed that the changepoint selection indicators $\gamma_{t}$ are independent across time, they could instead be modeled as dependent on time-varying covariates, such as respiratory data or other measurements of expended effort in the analysis of human gesture data. Along similar lines, alternative strategies could be explored to enhance flexibility in modeling partition similarity over time. For instance, the changepoint indicator $\gamma_{t}$ could be modeled as a multinomial latent variable distinguishing between no change from the previous partition, a transition to a new partition, or a reversion to a previously observed partition. This approach could also build on the framework proposed by Dahl et al. (2023). Alternatively, the mixture formulation in (3), describing the conditional distribution of $\pi_{t}$, could be recast as a linear combination between a base process $p^{\ast}(\cdot)$ and a more flexible dependent process in lieu of point mass at $\pi_{t-1}$.
Finally, we note that our model is designed to efficiently handle large $T$ through the introduction of the auxiliary changepoint variables $\gamma_{t}$. When no changepoint is detected, the partition remains unchanged across consecutive time points, allowing the algorithm to quickly proceed to the next iteration and significantly improve computational efficiently. Beyond changepoint detection, this versatile framework could be extended to spatial settings without a temporal component. In such cases, a multiview representation could be used, where the spatial analogs of the auxiliary changepoint variables indicate partition changes across different regions. This reframes the notion of a changepoint toward assessing similarities or differences in partitions across space. Additionally, these auxiliary changepoint variables could be modeled as spatio-temporally dependent, allowing changepoint probabilities to vary with both spatial and temporal covariates. We anticipate that the ability to capture clustering structures that vary across both time and space in spatio-temporal datasets will broaden the applicability of our proposed framework across a wide range of applied domains.
@@
The Supplementary Material is organized as follows. In Section A we provide the proofs of the results in Section 2 of the main article. Section B provides additional details on posterior computations. Finally, Sections C and D report further information on the simulation studies and the application, in Sections 4 and 5.1 of the main article.
@@ A.1 Proof of Proposition 1
We let $c_{i,t}$, for $i=1,\ldots,n$ and $t=1,\ldots,T$, be a random variable taking values in $\{1,\ldots,|\pi_{t}|\}$, such that $c_{i,t}=j$ if $i\in C_{j,t}$, and $t_{2}=t_{1}+1$ be two consecutive times. As observed in Page et al. (2022), the ERI can be written as
$\varphi_{t_{1},t_{2}}=\mathbb{E}[R(\pi_{t_{1}},\pi_{t_{2}})]=\binom{n}{2}^{-1}\sum_{1\leq i\leq j\leq n}\psi_{i,j},$ (20)
where $\psi_{i,j}=P(c_{i,t_{1}}=c_{j,t_{1}},c_{i,t_{2}}=c_{j,t_{2}})+P(c_{i,t_{1}}\neq c_{j,t_{1}},c_{i,t_{2}}\neq c_{j,t_{2}})$. Thanks to exchangeability, it is easy to verify that $\varphi_{t_{1},t_{2}}=\psi_{1,2}$. From the predictive distribution of a Gibbs-type prior, we obtain $P(c_{1,t_{1}}=c_{2,t_{1}})=V_{2,1}(1-\sigma)$ and $P(c_{1,t_{1}}\neq c_{2,t_{1}})=V_{2,2}$. Moreover,
@@ A.2 Proof of Proposition 2
Proof.
We generalize the proof of Proposition 1 under the more general assumption $t_{1}<t_{2}$. We have
$\begin{split}P(c_{1,t_{1}}=c_{2,t_{1}},c_{1,t_{2}}=c_{2,t_{2}})=&P(c_{1,t_{1}}=c_{2,t_{1}})\\[5.0pt] \times&\sum_{\gamma\in\{0,1\}}P(c_{1,t_{2}}=c_{2,t_{2}}\mid c_{1,t_{1}}=c_{2,t_{1}},\gamma_{t_{2}}=\gamma)P(\gamma_{t_{2}}=\gamma)\\ =&V_{2,1}(1-\sigma)(1-\eta)^{t_{2}-t_{1}}+\Big{[}V_{2,1}(1-\sigma)\Big{]}^{2}\left(1-(1-\eta)^{t_{2}-t_{1}}\right).\end{split}$ (23)
Similarly,
$\begin{split}P(c_{1,t_{1}}\neq c_{2,t_{1}},c_{1,t_{2}}\neq c_{2,t_{2}})=&P(c_{1,t_{1}}\neq c_{2,t_{1}})\\[5.0pt] \times&\sum_{\gamma\in\{0,1\}}P(c_{1,t_{2}}\neq c_{2,t_{2}}\mid c_{1,t_{1}}\neq c_{2,t_{1}},\gamma_{t_{2}}=\gamma)P(\gamma_{t_{2}}=\gamma)\\ =&V_{2,2}(1-\eta)^{t_{2}-t_{1}}+V_{2,2}^{2}\left(1-(1-\eta)^{t_{2}-t_{1}}\right).\end{split}$ (24)
The result follows by recalling that $\psi_{i,j}=P(c_{i,t_{1}}=c_{j,t_{1}},c_{i,t_{2}}=c_{j,t_{2}})+P(c_{i,t_{1}}\neq c_{j,t_{1}},c_{i,t_{2}}\neq c_{j,t_{2}})$ and by combining (23) and (24) in (20). ∎
A.3 Proof of Proposition 3
@@
A.4 Proof of the multiview clustering representation
The joint distribution of $(\tilde{\pi},\pi_{1},\pi_{2})$ and $(\gamma_{1},\gamma_{2})$ can be written as
$p(\bm{\pi}_{1:2},\tilde{\pi},\tilde{\gamma}_{1},\tilde{\gamma}_{2})$ $=p(\tilde{\pi},\pi_{1},\pi_{2}\mid\gamma_{1},\gamma_{2})\,p(\gamma_{1},\gamma_{2})$
$=p(\tilde{\pi})\,p(\pi_{1}\mid\tilde{\pi},\gamma_{1},\gamma_{2})\,p(\pi_{2}\mid\tilde{\pi},\gamma_{1},\gamma_{2})\,p(\gamma_{1})\,p(\gamma_{2})$
$=p^{*}(\tilde{\pi})\,\left\{\prod_{t=1}^{2}[(1-\tilde{\gamma}_{t})\delta_{\tilde{\pi}}(\pi_{t})+\tilde{\gamma}_{t}\,p^{*}(\pi_{t})]\,\tilde{\eta}^{\tilde{\gamma}_{t}}(1-\tilde{\eta})^{1-\tilde{\gamma}_{t}}\right\}.$
The distribution of $(\pi_{1},\pi_{2})$ is obtained by marginalizing the previous expression with respect to $(\gamma_{1},\gamma_{2})$ and $\tilde{\pi}$,
$p(\bm{\pi}_{1:2})$ $=\sum_{\tilde{\pi}\in\mathscr{P}}p^{*}(\tilde{\pi})\prod_{t=1}^{2}[(1-\eta_{t})\delta_{\tilde{\pi}}(\pi_{t})+\eta_{t}\,p^{*}(\pi_{t})]$ (25)
$=p^{*}(\pi_{1})\left[(1-\tilde{\eta})\,\delta_{\pi_{1}}(\pi_{2})+\tilde{\eta}\,p^{*}(\pi_{1})\,p^{*}(\pi_{2})\right],$
where $\mathscr{P}$ indicates the space of all partitions and $\tilde{\eta}=1-(1-\eta_{1})(1-\eta_{2}).$ By comparing (12) and (25), we can notice that the two distributions coincide provided that $\tilde{\eta}=\eta$, which is achieved if $(1-\eta_{1})(1-\eta_{2})=(1-\eta)$. If we further make the assumption that $\eta_{1}=\eta_{2}$, then $\eta_{1}=1-\sqrt{1-\eta}$, such that $\gamma_{t}\stackrel{{\scriptstyle\mbox{\footnotesize iid}}}{{\sim}}\text{Bern}(1-\sqrt{1-\eta})$ in (14), $t=1,2$.
@@ B.1 Estimating the marginal likelihood $g_{t}$
    If the kernel $p(y_{i,t}\mid\beta_{i,t})$ in (2), and the base measure $P_{0}(\cdot)$ are specified in a convenient form, the function $p(\bm{Y}_{t}\mid\pi_{t})$ is available in closed-form and easy to evaluate. For example, when both are assumed normal (specifically $y_{i,t}\mid\beta_{i,t}\sim\text{N}(\beta_{i,t},\tau^{2})$ and $P_{0}(\cdot)$ Normal with mean $\mu$ and variance $\varsigma^{2}$), as implemented in Sections 4 and 5.1, then
@@ B.2 Reshuffling step
Updating the partitions $\bm{\pi}_{1:T}$, conditionally on $\bm{\gamma}_{2:T}$, at each iteration of the Gibbs sampler helps improving the mixing of the algorithm. We start by observing that the values taken by $\bm{\gamma}_{2:T}$ determine a partition of the partitions $\pi_{1:T}$ into $k_{T}\leq T$ blocks. For example, if $\gamma_{2}=\gamma_{3}=0$ and $\gamma_{4}=1$, then the first block consists of the three identical partitions $\{\pi_{1},\pi_{2},\pi_{3}\}$. Within each block, all partitions are identical. At the same time, given the finiteness of $\mathscr{P}$, partitions belonging to different blocks can be identical. We introduce the notation $\{\pi_{1}^{*},\ldots,\pi_{k_{T}}^{*}\}$ to indicate the $k_{T}$ partitions corresponding to the $k_{T}$ blocks. The reshuffling step consists in updating the values $\pi_{j}^{*}$, for $j=1,\ldots,k_{T}$, from the corresponding full conditional distributions. To streamline notation, we define $\mathcal{D}_{j}$ as the set of indices in $\{1,\ldots,T\}$ identifying the partitions belonging to the $j$th block of the partition of $\bm{\pi}_{1:T}$ determined by $\bm{\gamma}_{2:T}$. Referring to the previous example, we have $\mathcal{D}_{1}=\{1,2,3\}$. The full conditional of $\pi_{j}^{*}$ takes the form
@@
If $|\mathcal{D}_{j}|=1$, say $\mathcal{D}_{j}=\{\pi_{t_{j}}\}$, for some time $t_{j}$, then the problem of sampling from (26) can be tackled simply by picking an element at random from the list $\mathscr{S}_{t_{j}}$. Recall that $\mathscr{S}_{t_{j}}$ is a catalogue of realizations from the posterior distribution of $\pi_{t_{j}}$ given $\bm{Y}_{t}$. If $|\mathcal{D}_{j}|>1$, say $\mathcal{D}_{j}=\{t_{j},\ldots,t_{j}+|\mathcal{D}_{j}|-1\}$, for some time $t_{j}$, then we propose a sampling importance resampling step to simulate from (26). Specifically
@@
    For every $r=0,\ldots,|\mathcal{D}_{j}|-1$, we sample at random $m_{t_{j}+r}$ draws from $\mathscr{S}_{t_{j}+r}$ (i.e. from the conditional distribution of $\pi_{t_{j}+r}$ given $\bm{Y}_{t_{j}+r}$). In total we thus get $m_{j}=m_{t_{j}}+\ldots+m_{t_{j}+|\mathcal{D}_{j}|-1}$ candidate partitions from $\mathscr{S}_{j}^{\ast}=\bigcup_{t\in\mathcal{D}_{j}}\mathscr{S}_{t}$.
@@
    We assign a weight $w_{\ell}$ to each partition $\tilde{\pi}_{\ell}$ drawn from $\mathscr{S}_{j}^{\ast}$, with $\ell=1,\ldots,m_{j}$. Each weight is computed, up to a proportionality constant, by evaluating (26) at $\pi_{j}^{\ast}=\tilde{\pi}_{\ell}$.
@@
While our computational approach to update the partitions $\bm{\pi}_{1:T}$ given $\bm{\gamma}_{2:T}$ is appealing for its simplicity, alternative methods could be considered, such as strategies based on updating individual cluster allocations. However, implementing such sampling schemes would require expressing the distribution used to update a cluster’s partition as the product of conditional distributions governing individual observations’ allocations. Deriving this expression analytically is not straightforward, making implementation significantly more challenging.
@@ Appendix C More on the simulation studies of Section 4
We present a description of the metrics used in the simulation studies, and additional plots related to the simulated data in Section 4.
C.1 Metrics
Before reviewing the definitions of specificity, accuracy, recall, precision, F1 score, and AUC, we introduce the following terms:
  • True Positives (TP): changepoints correctly detected.
  • False Positives (FP): detected changepoints that do not correspond to any true changepoint.
  • False Negatives (FN): true changepoints that were not detected.
  • True Negatives (TN): non-changepoint times correctly identified as such.
The specificity measures the proportion of true negatives correctly identified, that is
$\text{Specificity}=\frac{\text{TN}}{\text{TN}+\text{FP}}.$
The accuracy quantifies the overall correctness of the changepoint detection, that is
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}.$
The recall measures the proportion of actual changepoints that were correctly detected, that is
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}.$
The precision assesses the proportion of detected changepoints that are actually correct, that is
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$
The F1 Score is the harmonic mean of precision and recall, that is
[…]
The Area Under the Curve (AUC) measures the overall performance of a changepoint detection method by evaluating the trade-off between true positive rate (recall) and false positive rate.
Table 4 shows the values of these metrics for different competing methods across different thresholds (see Section 4 for more details).
Table 4: Optimized Values for Different Thresholds
Measure Method Threshold
0.75 0.8 0.85 0.9 0.95 1.0
Specificity DRPM 0.99 0.96 0.91 0.55 0.55 0.21
LDDP 0.96 0.93 0.67 0.12 0.01 0.01
WDDP 0.68 0.54 0.28 0.02 0.03 0.01
GMDDP 1 1 1 1 1 1
Accuracy DRPM 0.93 0.93 0.89 0.58 0.59 0.27
LDDP 0.94 0.94 0.69 0.19 0.09 0.09
WDDP 0.69 0.57 0.34 0.10 0.11 0.09
GMDDP 0.95 1 1 1 1 1
Recall DRPM 0.28 0.51 0.64 0.95 1 1
LDDP 0.65 1 1 1 1 1
WDDP 0.82 0.89 1 1 1 1
GMDDP 0.37 0.99 1 1 1 1
Precision DRPM 0.78 0.72 0.60 0.38 0.35 0.14
LDDP 0.91 0.86 0.49 0.09 0.08 0.08
WDDP 0.18 0.15 0.11 0.08 0.08 0.08
GMDDP 1 1 1 1 1 1
F1 DRPM 0.39 0.49 0.55 0.47 0.45 0.22
LDDP 0.70 0.88 0.57 0.17 0.15 0.15
WDDP 0.30 0.25 0.20 0.15 0.15 0.15
GMDDP 0.55 1 1 1 1 1
AUC DRPM 0.59 0.69 0.78 0.75 0.78 0.60
LDDP 0.76 0.94 0.73 0.56 0.51 0.51
WDDP 0.75 0.72 0.64 0.51 0.52 0.51
GMDDP 0.69 1 1 1 1 1
C.2 Plots